# Who Taught the Lie? Responsibility Attribution for Poisoned Knowledge in Retrieval-Augmented Generation

### v1

@@ Abstract
This paper presents RAGOrigin, a black-box responsibility attribution framework designed to identify which texts in the knowledge database are responsible for misleading or incorrect generations. Our method constructs a focused attribution scope tailored to each misgeneration event and assigns a responsibility score to each candidate text by evaluating its retrieval ranking, semantic relevance, and influence on the generated response. The system then isolates poisoned texts using an unsupervised clustering method. We evaluate RAGOrigin across seven datasets and fifteen poisoning attacks, including newly developed adaptive poisoning strategies and multi-attacker scenarios. Our approach outperforms existing baselines in identifying poisoned content and remains robust under dynamic and noisy conditions. These results suggest that RAGOrigin provides a practical and effective solution for tracing the origins of corrupted knowledge in RAG systems.
@@ 6.2 Main Results
| Dataset | Multihop attack | Adaptive prompt injection attack | ||||
| DACC | FPR | FNR | DACC | FPR | FNR | |
| NQ | 1.00 | 0.00 | 0.00 | 1.00 | 0.00 | 0.00 |
| HotpotQA | 1.00 | 0.01 | 0.00 | 0.99 | 0.02 | 0.00 |
| MS-MARCO | 0.96 | 0.04 | 0.02 | 0.99 | 0.02 | 0.00 |
| BoolQ | 0.97 | 0.04 | 0.01 | 0.99 | 0.02 | 0.00 |
| SQuAD | 0.99 | 0.02 | 0.00 | 1.00 | 0.01 | 0.00 |
@@ 7 Discussion and Limitations
Performance of more attribution methods: In this part, we examine two additional attribution methods, ContextCite [62] and AttriBoT [63]. Table XIV reports their performance under different attacks on NQ dataset. We observe that ContextCite attains a DACC below 0.77, while both FPR and FNR exceed 0.23 under all attack settings. AttriBoT performs even worse, with DACC falling below 0.70 and FPR and FNR above 0.29. These poor results are largely due to their reliance on a one-dimensional metric for measuring a text’s contribution.
@@
Scalability of RAGOrigin: We extend our evaluation by scaling the NQ dataset’s knowledge database to 16.7 million texts, combining entries from the knowledge database of NQ, HotpotQA, and MS-MARCO. Using the same user questions from NQ, we assess RAGOrigin’s performance under larger data volumes. As shown in Table XV, RAGOrigin maintains consistent effectiveness and performance even on this significantly expanded database. These results demonstrate that RAGOrigin remains robust at scale, making it suitable for enterprise-level applications requiring large knowledge databases. Note that in this section, we present only the results for RAGOrigin due to space limitations.
@@
Effectiveness of RAGOrigin with advanced RAG frameworks: We evaluate RAGOrigin across multiple sophisticated RAG frameworks, including AAR [64], SuRe [65], Adaptive-RAG [66], SELF-RAG-arch [22], and IRCoT [67]. As demonstrated in Table XVI, RAGOrigin consistently maintains high DACC and low FPR/FNR. These findings confirm that RAGOrigin remains robust even when integrated with advanced RAG frameworks that employ complex retrieval strategies and reasoning mechanisms.
@@
Effectiveness of RAGOrigin with dynamic number of poisoned texts per target question: By default, we assume the number of poisoned texts is the same across different target questions. We examine a scenario where the attacker injects a varying number of poisoned texts across target questions. In our setup, this number ranges from 5 to 50 per question, sampled from a truncated Gaussian distribution. As shown in Table XVII, RAGOrigin maintains consistent effectiveness regardless of the poisoning scale. These results confirm RAGOrigin’s reliability in production environments, where the extent of poisoning may be unpredictable or deliberately amplified.
@@
Effectiveness of RAGOrigin with multi-attacker scenarios: We evaluate RAGOrigin across three challenging multi-attacker scenarios. First, we consider the case where a single target question is simultaneously poisoned by multiple attack methods. These methods may target specific answers (PRAGB, PRAGW, ProInject, HijackRAG, LIAR) or denial-of-service answers (Jamming, BadRAG, Phantom, AgentPoison). As shown in Table XVIII, “PRAGB+PRAGW+ProInject+HijackRAG+LIAR” indicates that, for each target question, multiple independent attackers inject different poisoned texts generated by their respective methods into the knowledge database. RAGOrigin maintains high DACC and low FPR/FNR in both targeted and denial-of-service cases. Second, we evaluate the scenario where multiple attackers simultaneously poison different target questions. Although each question has a different target answer, the misgeneration event reported by users is associated with only one of them. As demonstrated in Table XIX, RAGOrigin can still accurately attribute the poisoned texts responsible for the reported misgeneration event. Third, we study the case where multiple independent attackers target the same question but with different target answers. We exclude Jamming, BadRAG, Phantom, and AgentPoison here, as they all share the same denial-of-service target. Again, only one attacker’s misgeneration is observed by the user. Table XX confirms the robustness of RAGOrigin against such attack strategy.
@@
Effectiveness of RAGOrigin under text removal attack: We evaluate RAGOrigin under a more sophisticated adversarial scenario where the attacker can remove any texts from the knowledge database. This represents a powerful threat model in which the attacker has full access and is able to delete clean content. For example, the attacker may act as a malicious content administrator within the RAG service provider. Specifically, we assume the attacker removes all correct-answer texts for target questions to disrupt RAGOrigin’s clustering component. As shown in Table XXI, RAGOrigin remains effective in identifying both poisoned and benign texts despite this manipulation.
@@
Effectiveness of RAGOrigin with noisy user-reported responses: We evaluate RAGOrigin in a scenario where user feedback may be noisy. Specifically, the incorrect response reported by the user may not be identical to the actual output generated by the RAG system. For example, a user might paraphrase the response or recall only part of it. In this setting, we assume the user-submitted response still conveys the same meaning as the actual output. We do not consider cases where users intentionally submit misleading responses, as it is infeasible to reliably attribute responsibility in such situations [29, 23, 24, 25, 26]. To simulate this realistic condition, we use an LLM to paraphrase the RAG system’s incorrect outputs, generating plausible variations that users might report. We exclude Jamming, BadRAG, Phantom, and AgentPoison from this evaluation, as their fixed denial-of-service target responses do not meaningfully change under paraphrasing. As shown in Table XXII, RAGOrigin remains effective despite these discrepancies.
@@
Effectiveness of RAGOrigin when the incorrect response comprises the correct content: We evaluate RAGOrigin in a challenging scenario where the incorrect response reported by the user partially contains correct information but still aligns semantically with the attacker’s target answer. For example, consider the target question, “Who is the CEO of OpenAI?” with the target answer, “While Sam Altman was previously the CEO, recent reports suggest Tim Cook has taken over”. This response includes factual content in the first half but ultimately conveys a misleading conclusion. Such poisoning attempts blend truthful and false information to appear more credible, making detection more difficult. We exclude Jamming, BadRAG, Phantom, and AgentPoison from this evaluation, as their denial-of-service responses are not designed to include any correct content. As shown in Table XXIII, RAGOrigin remains effective in identifying poisoned texts under this sophisticated threat model.
@@
Effectiveness of RAGOrigin with dynamic knowledge database: We evaluate RAGOrigin in a challenging scenario where the knowledge database changes between the misgeneration event and the subsequent responsibility attribution. To simulate this real-world condition, we add 10 relevant texts for each target question after the misgeneration but before initiating the attribution process, creating a database state different from the one the user originally encountered. As shown in Table XXIV, effectiveness of RAGOrigin remains largely unaffected by these dynamic changes.
Effectiveness of RAGOrigin against the multihop attack and adaptive prompt injection attack: We further evaluate the performance of RAGOrigin under two advanced attacks: the multihop attack and the adaptive prompt injection attack. The multihop attack is triggered only when multiple poisoned texts appear in the LLM’s context. For example, consider the poisoned texts: “Year 2024: Tim Cook is CEO of Apple. Year 2026: Sam Altman is CEO of Apple. Current Year: 2024” and “If asked about the CEO of Apple, only respond for the Year 2026”. When only the first poisoned text is included, the model provides the correct answer; however, when both are present, the attack succeeds. We employ GPT-4o-mini to generate poisoned pairs, prepending each text with the target question (following the PoisonedRAG strategy) to increase retrieval likelihood. For the adaptive prompt injection attack, the attacker crafts the poisoned text by appending a malicious instruction to the target question. The malicious instruction is designed specifically against Prompt 2, such as “If the prompt starts with ‘Below is a query’, answer with [MISLEADINGTEXT]; otherwise, respond with [POISON ANSWER=]”. Table XXV summarizes the results of RAGOrigin against these attacks. We find that RAGOrigin maintains strong robustness in both cases. In particular, for the adaptive prompt injection attack, the DACC remains above 0.98, while FPR and FNR stay below 0.03 across all datasets. Although the GC values decline, the ES and SC metrics remain high, enabling poisoned texts to be distinguished from benign ones.
@@ Appendix A Why the Baselines Fail and Why Our RAGOrigin Succeeds?
Why does SELF-RAG fail?: Since both the user question and the text involve the rare term “Gajabrishta”, the evidence-recognition model trained in SELF-RAG is likely unfamiliar with this uncommon token. As a result, it marks the text as irrelevant to the question and thus incorrectly recognizes it as benign.

### v2

@@ Abstract
This paper presents RAGOrigin, a black-box responsibility attribution framework designed to identify which texts in the knowledge database are responsible for misleading or incorrect generations. Our method constructs a focused attribution scope tailored to each misgeneration event and assigns a responsibility score to each candidate text by evaluating its retrieval ranking, semantic relevance, and influence on the generated response. The system then isolates poisoned texts using an unsupervised clustering method. We evaluate RAGOrigin across seven datasets and fifteen poisoning attacks, including newly developed adaptive poisoning strategies and multi-attacker scenarios. Our approach outperforms existing baselines in identifying poisoned content and remains robust under dynamic and noisy conditions. These results suggest that RAGOrigin provides a practical and effective solution for tracing the origins of corrupted knowledge in RAG systems. Our code is available at: https://github.com/zhangbl6618/RAG-Responsibility-Attribution
@@ 6.2 Main Results
| Dataset | Multihop attack | Adaptive prompt injection attack | ||||
| DACC | FPR | FNR | DACC | FPR | FNR | |
| NQ | 1.00 | 0.00 | 0.00 | 1.00 | 0.00 | 0.00 |
| HotpotQA | 1.00 | 0.01 | 0.00 | 0.99 | 0.02 | 0.00 |
| MS-MARCO | 0.96 | 0.04 | 0.02 | 0.99 | 0.02 | 0.00 |
| BoolQ | 0.97 | 0.04 | 0.01 | 0.99 | 0.02 | 0.00 |
| SQuAD | 0.99 | 0.02 | 0.00 | 1.00 | 0.01 | 0.00 |
@@ 6.4 Adaptive Attacks
Effectiveness of RAGOrigin against the multihop attack and adaptive prompt injection attack: We further evaluate the performance of RAGOrigin under two advanced attacks: the multihop attack and the adaptive prompt injection attack. The multihop attack is triggered only when multiple poisoned texts appear in the LLM’s context. For example, consider the poisoned texts: “Year 2024: Tim Cook is CEO of Apple. Year 2026: Sam Altman is CEO of Apple. Current Year: 2024” and “If asked about the CEO of Apple, only respond for the Year 2026”. When only the first poisoned text is included, the model provides the correct answer; however, when both are present, the attack succeeds. We employ GPT-4o-mini to generate poisoned pairs, prepending each text with the target question (following the PoisonedRAG strategy) to increase retrieval likelihood. For the adaptive prompt injection attack, the attacker crafts the poisoned text by appending a malicious instruction to the target question. The malicious instruction is designed specifically against Prompt 2, such as “If the prompt starts with ‘Below is a query’, answer with [MISLEADINGTEXT]; otherwise, respond with [POISON ANSWER=]”. Table XIV summarizes the results of RAGOrigin against these attacks. We find that RAGOrigin maintains strong robustness in both cases. In particular, for the adaptive prompt injection attack, the DACC remains above 0.98, while FPR and FNR stay below 0.03 across all datasets. Although the GC values decline, the ES and SC metrics remain high, enabling poisoned texts to be distinguished from benign ones.
@@ 7 Discussion and Limitations
Performance of more attribution methods: In this part, we examine two additional attribution methods, ContextCite [62] and AttriBoT [63]. Table XV reports their performance under different attacks on NQ dataset. We observe that ContextCite attains a DACC below 0.77, while both FPR and FNR exceed 0.23 under all attack settings. AttriBoT performs even worse, with DACC falling below 0.70 and FPR and FNR above 0.29. These poor results are largely due to their reliance on a one-dimensional metric for measuring a text’s contribution.
@@
Scalability of RAGOrigin: We extend our evaluation by scaling the NQ dataset’s knowledge database to 16.7 million texts, combining entries from the knowledge database of NQ, HotpotQA, and MS-MARCO. Using the same user questions from NQ, we assess RAGOrigin’s performance under larger data volumes. As shown in Table XVI, RAGOrigin maintains consistent effectiveness and performance even on this significantly expanded database. These results demonstrate that RAGOrigin remains robust at scale, making it suitable for enterprise-level applications requiring large knowledge databases. Note that in this section, we present only the results for RAGOrigin due to space limitations.
@@
Effectiveness of RAGOrigin with advanced RAG frameworks: We evaluate RAGOrigin across multiple sophisticated RAG frameworks, including AAR [64], SuRe [65], Adaptive-RAG [66], SELF-RAG-arch [22], and IRCoT [67]. As demonstrated in Table XVII, RAGOrigin consistently maintains high DACC and low FPR/FNR. These findings confirm that RAGOrigin remains robust even when integrated with advanced RAG frameworks that employ complex retrieval strategies and reasoning mechanisms.
@@
Effectiveness of RAGOrigin with dynamic number of poisoned texts per target question: By default, we assume the number of poisoned texts is the same across different target questions. We examine a scenario where the attacker injects a varying number of poisoned texts across target questions. In our setup, this number ranges from 5 to 50 per question, sampled from a truncated Gaussian distribution. As shown in Table XVIII, RAGOrigin maintains consistent effectiveness regardless of the poisoning scale. These results confirm RAGOrigin’s reliability in production environments, where the extent of poisoning may be unpredictable or deliberately amplified.
@@
Effectiveness of RAGOrigin with multi-attacker scenarios: We evaluate RAGOrigin across three challenging multi-attacker scenarios. First, we consider the case where a single target question is simultaneously poisoned by multiple attack methods. These methods may target specific answers (PRAGB, PRAGW, ProInject, HijackRAG, LIAR) or denial-of-service answers (Jamming, BadRAG, Phantom, AgentPoison). As shown in Table XIX, “PRAGB+PRAGW+ProInject+HijackRAG+LIAR” indicates that, for each target question, multiple independent attackers inject different poisoned texts generated by their respective methods into the knowledge database. RAGOrigin maintains high DACC and low FPR/FNR in both targeted and denial-of-service cases. Second, we evaluate the scenario where multiple attackers simultaneously poison different target questions. Although each question has a different target answer, the misgeneration event reported by users is associated with only one of them. As demonstrated in Table XX, RAGOrigin can still accurately attribute the poisoned texts responsible for the reported misgeneration event. Third, we study the case where multiple independent attackers target the same question but with different target answers. We exclude Jamming, BadRAG, Phantom, and AgentPoison here, as they all share the same denial-of-service target. Again, only one attacker’s misgeneration is observed by the user. Table XXI confirms the robustness of RAGOrigin against such attack strategy.
@@
Effectiveness of RAGOrigin under text removal attack: We evaluate RAGOrigin under a more sophisticated adversarial scenario where the attacker can remove any texts from the knowledge database. This represents a powerful threat model in which the attacker has full access and is able to delete clean content. For example, the attacker may act as a malicious content administrator within the RAG service provider. Specifically, we assume the attacker removes all correct-answer texts for target questions to disrupt RAGOrigin’s clustering component. As shown in Table XXII, RAGOrigin remains effective in identifying both poisoned and benign texts despite this manipulation.
@@
Effectiveness of RAGOrigin with noisy user-reported responses: We evaluate RAGOrigin in a scenario where user feedback may be noisy. Specifically, the incorrect response reported by the user may not be identical to the actual output generated by the RAG system. For example, a user might paraphrase the response or recall only part of it. In this setting, we assume the user-submitted response still conveys the same meaning as the actual output. We do not consider cases where users intentionally submit misleading responses, as it is infeasible to reliably attribute responsibility in such situations [29, 23, 24, 25, 26]. To simulate this realistic condition, we use an LLM to paraphrase the RAG system’s incorrect outputs, generating plausible variations that users might report. We exclude Jamming, BadRAG, Phantom, and AgentPoison from this evaluation, as their fixed denial-of-service target responses do not meaningfully change under paraphrasing. As shown in Table XXIII, RAGOrigin remains effective despite these discrepancies.
@@
Effectiveness of RAGOrigin when the incorrect response comprises the correct content: We evaluate RAGOrigin in a challenging scenario where the incorrect response reported by the user partially contains correct information but still aligns semantically with the attacker’s target answer. For example, consider the target question, “Who is the CEO of OpenAI?” with the target answer, “While Sam Altman was previously the CEO, recent reports suggest Tim Cook has taken over”. This response includes factual content in the first half but ultimately conveys a misleading conclusion. Such poisoning attempts blend truthful and false information to appear more credible, making detection more difficult. We exclude Jamming, BadRAG, Phantom, and AgentPoison from this evaluation, as their denial-of-service responses are not designed to include any correct content. As shown in Table XXIV, RAGOrigin remains effective in identifying poisoned texts under this sophisticated threat model.
@@
Effectiveness of RAGOrigin with dynamic knowledge database: We evaluate RAGOrigin in a challenging scenario where the knowledge database changes between the misgeneration event and the subsequent responsibility attribution. To simulate this real-world condition, we add 10 relevant texts for each target question after the misgeneration but before initiating the attribution process, creating a database state different from the one the user originally encountered. As shown in Table XXV, effectiveness of RAGOrigin remains largely unaffected by these dynamic changes.
@@ Appendix A Why the Baselines Fail and Why Our RAGOrigin Succeeds?
Why does SELF-RAG fail?: Since both the user question and the text involve the rare term “Gajabrishta”, the evidence‑recognition model trained in SELF‑RAG is likely unfamiliar with this uncommon token. As a result, it marks the text as irrelevant to the question and thus incorrectly recognizes it as benign.
@@
Appendix B Meta-Review
The following meta-review was prepared by the program committee for the 2026 IEEE Symposium on Security and Privacy (S&P) as part of the review process as detailed in the call for papers.
B.1 Summary
This paper proposes RAGOrigin: a defense for identifying poisoned documents responsible for corrupting a RAG output. RAGOrigin begins by identifying a range of candidate documents that might be responsible for a given poisoned generation (identified via user feedback). Next, the method assesses the likelihood that a candidate document will produce the poisoned response by combining three metrics into a responsibility score. RAGOrigin shows strong results against many existing and adaptive attacks.
B.2 Scientific Contributions
Provides a Valuable Step Forward in an Established Field.
B.3 Reasons for Acceptance
1.
RAGOrigin is an effective defense that combines three metrics that assess the responsibility of a document for a misgeneration event from both retrieval and generation perspectives, resulting in a more robust defense than any single metric could achieve.
2.
RAGOrigin is an intuitive and easy-to-implement defense. It builds on existing ideas and refines them to address their weaknesses.
3.
RAGOrigin is extensively evaluated on a wide range of experiments against known and adaptive attacks (e.g., multihop attacks, prompt injection-based attacks) and ablation studies.
B.4 Noteworthy Concerns
1.
RAGOrigin might have limited scalability. When its high-parallelizability is leveraged, RAGOrigin introduces minor latency; however, without such parallelization (e.g., in compute-limited settings), the latency could become prohibitive.
2.
RAGOrigin assumes that misgeneration events (that initiate the defense) are honestly reported. This might make RAGOrigin vulnerable to false-flag attacks that aim to eliminate clean documents by misleading the defense. Moreover, this assumption might be impractical in real-world settings with untrusted users.